# Finding the Fairest Voting System Using Likelihood Analysis

### v2

@@ A.1 IIA
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
Table 11: Percentage IIA Violations Minimax
@@
3 100.00 100.00 100.00 100.00
4 100.00 100.00 100.00 100.00
5 100.00 100.00 100.00 100.00
10 100.00 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 99.99 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 99.99 100.00 100.00 100.00
Table 12: Percentage IIA Violations Borda
@@
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.90 100.00 100.00
30 84.52 99.87 100.00 100.00
50 84.54 99.90 100.00 100.00
100 84.14 99.93 100.00 100.00
1000 83.64 99.91 100.00 100.00
5000 83.60 99.87 100.00 100.00
10000 84.12 99.90 100.00 100.00
30000 83.74 99.85 100.00 100.00
50000 83.96 99.76 100.00 100.00
Table 13: Percentage IIA Violations Coombs
@@
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.91 100.00 100.00
30 87.40 100.00 100.00 100.00
50 86.41 99.97 100.00 100.00
100 85.70 100.00 100.00 100.00
1000 85.52 99.99 100.00 100.00
5000 85.57 100.00 100.00 100.00
10000 85.91 99.99 100.00 100.00
30000 85.43 99.99 100.00 100.00
50000 85.45 99.98 100.00 100.00
Table 14: Percentage IIA Violations Dowdall
@@
3 89.08 99.08 99.96 100.00
4 95.03 99.81 100.00 100.00
5 78.62 99.99 100.00 100.00
10 91.80 100.00 100.00 100.00
30 96.72 100.00 100.00 100.00
50 97.29 100.00 100.00 100.00
100 96.91 100.00 100.00 100.00
1000 97.03 100.00 100.00 100.00
5000 96.66 100.00 100.00 100.00
10000 96.93 100.00 100.00 100.00
30000 96.83 99.99 100.00 100.00
50000 96.97 100.00 100.00 100.00
Table 15: Percentage IIA Violations Copeland
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 99.95 100.00 100.00 100.00
10 99.99 100.00 100.00 100.00
30 100.00 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 97.30 99.99 100.00 100.00
10 97.75 100.00 100.00 100.00
30 99.80 100.00 100.00 100.00
50 99.99 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 99.99 100.00 100.00 100.00
5000 99.99 100.00 100.00 100.00
10000 99.98 100.00 100.00 100.00
30000 99.97 100.00 100.00 100.00
50000 99.98 100.00 100.00 100.00
Table 16: Percentage IIA Violations Baldwin
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.35 99.78 99.98 100.00
4 73.63 99.96 100.00 100.00
5 92.87 99.39 100.00 100.00
10 94.79 99.99 100.00 100.00
30 99.96 100.00 100.00 100.00
50 100.00 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
1000 100.00 100.00 100.00 100.00
5000 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
30000 100.00 100.00 100.00 100.00
50000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.08 99.08 99.96 100.00
4 56.66 99.81 100.00 100.00
5 78.62 97.75 100.00 100.00
10 74.29 99.76 100.00 100.00
30 83.23 99.76 100.00 100.00
50 83.38 99.78 100.00 100.00
100 83.35 99.75 100.00 100.00
1000 82.49 99.76 100.00 100.00
5000 82.63 99.73 100.00 100.00
10000 83.14 99.74 100.00 100.00
30000 82.74 99.74 100.00 100.00
50000 82.87 99.70 100.00 100.00
A.2 Unanimity
Table 17: Percentage U Violations Plurality and Top Two
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 16.23 58.00 86.12 96.61
3 2.56 17.38 41.34 62.30
4 0.43 4.59 13.93 28.63
5 0.05 1.14 4.58 10.87
10 0.00 0.00 0.00 0.07
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 59.94 86.13 96.56
3 11.02 21.64 41.69 63.20
4 4.42 7.96 15.58 28.35
5 2.32 2.83 5.00 10.14
10 0.19 0.06 0.04 0.05
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
Table 18: Percentage U Violations Minimax
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 16.23 42.04 65.70 81.92
3 2.56 8.90 18.46 29.53
4 0.43 1.61 3.72 6.99
5 0.05 0.34 0.73 1.21
10 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
1 100.00 100.00 100.00 100.00
2 28.05 44.53 65.78 82.17
3 11.02 12.93 19.29 30.03
4 4.42 4.04 4.56 6.42
5 2.32 0.98 0.89 1.23
10 0.19 0.01 0.00 0.00
30 0.01 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
A.3 JS Frequencies
Recall that the JS Frequencies for Plurality, Instant Runoff, Minimax, and Top Two are always less than 0.01% in elections with at least three alternatives and two voters. For the rest of the systems, the JS Frequency tables are below.
Table 19: JS Frequencies Borda Count
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 5.16 0.01 0.00 0.00
30 0.11 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.12 0.00 0.00
30 15.48 0.07 0.00 0.00
50 15.71 0.18 0.00 0.00
100 16.15 0.06 0.00 0.00
1000 16.00 0.15 0.00 0.00
5000 15.48 0.14 0.00 0.00
10000 15.37 0.15 0.00 0.00
30000 15.89 0.09 0.00 0.00
50000 15.75 0.14 0.00 0.00
Table 20: JS Frequencies Coombs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 5.16 0.01 0.00 0.00
30 0.02 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.10 0.00 0.00
30 12.81 0.01 0.00 0.00
50 13.96 0.00 0.00 0.00
100 14.50 0.00 0.00 0.00
1000 14.20 0.02 0.00 0.00
5000 13.90 0.02 0.00 0.00
10000 13.40 0.01 0.00 0.00
30000 14.29 0.02 0.00 0.00
50000 14.02 0.01 0.00 0.00
Table 21: JS Frequencies Dowdall
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 0.22 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 7.41 0.02 0.00 0.00
30 3.30 0.00 0.00 0.00
50 3.27 0.00 0.00 0.00
100 3.16 0.00 0.00 0.00
1000 3.13 0.01 0.00 0.00
5000 3.11 0.00 0.00 0.00
10000 2.97 0.00 0.00 0.00
30000 2.98 0.00 0.00 0.00
50000 3.10 0.00 0.00 0.00
Table 22: JS Frequencies Copeland
(a) IC Data
Number of voters Number of alternatives
3       4       5 6
10 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3       4       5 6
10 1.90 0.00 0.00 0.00
30 0.15 0.00 0.00 0.00
50 0.08 0.00 0.00 0.00
100 0.02 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.02 0.00 0.00 0.00
10000 0.03 0.00 0.00 0.00
30000 0.01 0.00 0.00 0.00
50000 0.02 0.00 0.00 0.00
Table 23: JS Frequencies Baldwin
(a) IC Data
Number of voters Number of alternatives
3      4      5 6
10 5.16 0.01 0.00 0.00
30 0.10 0.00 0.00 0.00
50 0.00 0.00 0.00 0.00
100 0.00 0.00 0.00 0.00
1000 0.00 0.00 0.00 0.00
5000 0.00 0.00 0.00 0.00
10000 0.00 0.00 0.00 0.00
30000 0.00 0.00 0.00 0.00
50000 0.00 0.00 0.00 0.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 25.16 0.25 0.01 0.00
30 16.87 0.27 0.00 0.00
50 16.79 0.28 0.00 0.00
100 17.14 0.16 0.00 0.00
1000 17.27 0.22 0.00 0.00
5000 16.63 0.27 0.00 0.00
10000 16.68 0.29 0.00 0.00
30000 17.07 0.22 0.00 0.00
50000 16.88 0.33 0.00 0.00
Table 24: JS Frequencies Pairwise Majority
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
10 76.19 45.27 21.02 7.50
30 82.69 56.23 31.23 13.97
50 84.64 59.83 35.31 17.34
100 86.92 64.02 39.92 21.08
1000 90.05 70.89 48.37 28.82
5000 90.65 72.55 50.99 30.84
10000 90.81 72.64 51.11 31.40
30000 90.89 73.53 52.06 31.89
50000 91.08 73.77 52.04 32.39
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
10 84.50 51.12 22.47 7.69
30 90.42 64.15 33.53 14.42
50 91.81 67.86 38.56 17.78
100 92.77 71.49 43.87 21.85
1000 93.54 75.46 52.17 30.32
5000 93.76 75.97 53.24 32.27
10000 93.67 75.87 53.28 32.92
30000 93.55 75.94 53.08 33.02
50000 93.66 76.10 53.51 32.80

### v3

@@ A.3 Joint Violation Frequencies
Table 19: JV Frequencies Borda Count
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.30 99.87 100.00 100.00
4 73.42 100.00 100.00 100.00
10 94.84 99.99 100.00 100.00
30 99.89 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.38 99.17 99.92 100.00
4 55.94 99.81 99.99 100.00
10 74.84 99.88 100.00 100.00
30 84.52 99.93 100.00 100.00
100 83.85 99.94 100.00 100.00
10000 84.63 99.85 100.00 100.00
Table 20: JV Frequencies Coombs
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.30 99.87 100.00 100.00
4 73.42 100.00 100.00 100.00
10 94.84 99.99 100.00 100.00
30 99.98 100.00 100.00 100.00
100 100.00 100.00 100.00 100.00
10000 100.00 100.00 100.00 100.00
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.38 99.17 99.92 100.00
4 55.94 99.81 99.99 100.00
10 74.84 99.90 100.00 100.00
30 87.19 99.99 100.00 100.00
100 85.50 100.00 100.00 100.00
10000 86.60 99.99 100.00 100.00
Table 21: JV Frequencies Dowdall
(a) IC Data
Number of voters Number of alternatives
3 4 5 6
3 97.30 99.87 100.00 100.00
4 99.41 100.00 100.00 100.00
10 99.78 100.00 100.00 100.00
@@
(b) IAC Data
Number of voters Number of alternatives
3 4 5 6
3 89.38 99.17 99.92 100.00
4 94.94 99.81 99.99 100.00
10 92.59 99.98 100.00 100.00
30 96.70 100.00 100.00 100.00
100 96.84 100.00 100.00 100.00
10000 97.03 100.00 100.00 100.00
Table 22: JV Frequencies Copeland
@@
3 89.38 99.17 99.92 100.00
4 55.94 99.81 99.99 100.00
10 98.10 100.00 99.99 100.00
30 99.85 100.00 100.00 100.00
100 99.98 100.00 100.00 100.00
10000 99.97 100.00 100.00 100.00
Table 23: JV Frequencies Baldwin
@@
3 89.38 99.17 99.14 100.00
4 55.94 99.81 99.10 100.00
10 74.84 99.75 99.93 100.00
30 83.13 99.73 100.00 100.00
100 82.86 99.84 100.00 100.00
10000 83.32 99.71 100.00 100.00
Table 24: JV Frequencies Ranked Pairs
@@
3 49.20 89.57 99.14 99.97
4 55.94 89.50 99.10 99.97
10 72.34 97.72 99.93 100.00
30 74.57 98.29 100.00 100.00
100 75.66 98.67 100.00 100.00
10000 76.31 98.55 100.00 100.00
Table 25: JV Frequencies Pairwise Majority
@@
3 3.57 14.97 31.66 48.87
4 23.65 60.59 85.05 95.49
10 15.50 48.88 77.53 92.31
30 9.58 35.85 66.47 85.58
100 7.23 28.51 56.13 78.15
10000 6.34 24.13 46.72 67.08
Appendix B IIA Algorithm
The IIA algorithm used in this paper is different than that of Dougherty and Heckelman (2020) and we describe those differences here.
As IIA is an inter-profile condition, to check if a voting system violates IIA, multiple preference schedules have to be compared. Our IIA algorithm loops through all possible pairs of alternatives. For each alternative pair $A$ and $B$, the following test is done: First, find the societal preference between $A$ and $B$. Then, let $j$ equal the number of voters who prefer $A$ over $B$ and $k$ the number of voters who prefer $B$ over $A$. Then randomly generate 500 additional preference schedules in which $j$ voters still prefer $A$ over $B$ and $k$ voters prefer $B$ over $A$. The number 500 was chosen based on computational resources and run time. If any of these new preference schedules produce a different societal preference between $A$ and $B$ than the original, an IIA violation occurs. This is repeated for every possible pair of alternatives.
This algorithm can miss violations, which is more likely when the number of voters is high. As the number of voters increases, there are more possible preference schedules. Thus, in the 500 new preference schedules generated, there may be no preference schedule that has a different societal order, even if one exists. Checking every single preference schedule is impractical since there are $(\frac{c!}{2})^{v}$ possible schedules such that each voter maintains their relative preference between a pair, where $c$ is the number of alternatives and $v$ is the number of voters.
In Dougherty and Heckelman (2020), a different IIA algorithm was used. Our algorithm finds substantially more IIA violations, in particular with the IAC data. For Dowdall, Coombs, Copeland, and Baldwin our algorithm reports a greater percentage IIA violations. For Coombs and Copeland, the difference in the IAC data is around 20 percentage points.
B.0.1 Disagreement in IIA Conclusions under Copeland
2 0 0 1 0 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
Under Copeland, the societal preference order is $A\succ B\succ C$. Now, consider the pair of alternatives $A$ and $B$. Suppose that one voter who has preference $A\succ B\succ C$ changes their preference to $C\succ A\succ B$. Now the preference schedule is:
1 0 0 1 1 0
1st A A B B C C
2nd B C A C A B
3rd C B C A B A
Dougherty and Heckelman (2020) also state that Copeland violates IIA at a greater frequency for even number of voters $N$ than for the adjacent odd values N1N-1 and $N+1$ as there are more ties for elections with an even number of voters. Our data refutes this claim, showing that IIA violations in 3-alternative elections are greater for elections with 3 voters (97.35%) than with 4 voters (73.63%).